Question 1. (15 points) Suppose that Ken cares only about bathing suits (B) and flip-flops (F). His utility function is U = B 0.6F 0.4. The price of bathing suits are $10, and the price of flip-flops are $4. Ken has a budget of $100. (For Version 2, Income is $200 but prices remain the same).

(d) (5 points) What is the optimal choice of B and F for Ken to maximize his utility?
Answer #1

V= Bobo.4 PB $10 PF = $4 Budget = $100 MUB -0.4 F 0.4 0.6B At optimality MRS= mur = O.UBO GF-0.6 3 1/2 = 215 - 13F=5B] lo B+